OpenStudy (anonymous):

Find the nature of the series using raabe-duhamel criteria or ratio criteria(D'Alembert)

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\ln(1+a^n),a \ge0\]

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

I think all both ratio test and raabe are inconclusive you need to try something else

ganeshie8 (ganeshie8):

@eliassaab @SithsAndGiggles @dan815

OpenStudy (dan815):

so what are these criteria raabe-duhamel criteria or ratio criteria

OpenStudy (dan815):

i think there will be like a radius of convergence given some a

OpenStudy (dan815):

find that a

OpenStudy (dan815):

a<=1

OpenStudy (dan815):

but im pretty sure they want more than that, nature maybe HOW its converging and diverging as a function of a

OpenStudy (dan815):

for the sum itself u are dealing with a different a

OpenStudy (dan815):

like if a =1, u will have inf * ln (2)so diverging

OpenStudy (dan815):

u have to find the critical 'a'

OpenStudy (anonymous):

@dan815 it's pretty clear that this exercise must be done using one of those criteria

OpenStudy (dan815):

i think it comes down to rate of decrease of a^n when a<1 and rate of increase of ln function

OpenStudy (dan815):

oh okay then can u post that criteria stuff here

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

why isn't it \(a<1\)?

OpenStudy (anonymous):

ratio test seems to work, unless my algebra is bad

OpenStudy (anonymous):

Yes the ratio test work, using L'Hospital's rule. \[ \lim_{n->\infty}\frac {\ln(1+ a^{n+1})}{\ln(1+ a^{n)}}=\\ \lim_{n->\infty}\frac{a \left(a^n+1\right)}{a^{n+1}+1}=a \]

OpenStudy (anonymous):

Of course one has to suppose that a <1, since for \( a \ge 1 \), the series divereges

OpenStudy (anonymous):

I'm going to post my full solvation of this series

OpenStudy (anonymous):